assume an investor with the following utility function : U= E(r) - 0.60 (s2)

to maximize her expected utility, which one of the following investment alternatives would she choose?

A. a portfolio that pays 10% with a 60% probablility or 5% with 40% probability
B. A portfolio that pays 12% with 40% probability or 5% with 60% probability
C. A portfolio that pays 10% with 40% probability or 5% with a 60% probability
D. A portfolio that pays 12% with 60% probability or 5% with 40% probability
Answer #1

U = E(r) - 0.60(s2)

s is the standard deviation and E(r) is the expected return. So, s2 is the variance.

Please find the image of the excel.

A B с D E F 20 21 22 23 24 25 26 27 28 29 30 Probability Return 0.6 10% 0.4 5% Expected Return U Probability * Return Varianc

A B С D E F Probability Return Variance 0.6 0.4 0.1 0.05 Expected Return U Probability* Return =D25*C25 =D26*C26 =SUM(E25:E26

Thus, the correct asnwer is option D as U is maximised in D.

Comment in case of any query. Thumbs up would be highly appreciated.